An alloy containing 15% brass is to be combined with an alloy containing 35% brass to form an alloy containing 27% brass. How much of each alloy should be combined to make 200 pounds of the 27% brass alloy?
Answer by ankor@dixie-net.com(22740)

An alloy containing 15% brass is to be combined with an alloy containing 35% brass to form an alloy containing 27% brass.
How much of each alloy should be combined to make 200 pounds of the 27% brass alloy?
:
Let x = amt of 35% brass
the total is to be 200 lb, therefore
(200-x) = amt of 15% brass
:
.35x + .15(200-x) = .27(200)
.35x + 30 - .15x = 54
.35x - .15x = 54 - 30
.20x = 24
x = 24/.2
x = 120 lb of 35% brass
and
200 - 120 = 80 lb of 15% brass